# XAI-FUNGI: Dataset resulting from the user study on comprehensibility of explainable AI algorithms

### v1

@@ Abstract
1 Background & Summary
With the rapid development of black-box machine learning (ML) models, such as deep neural networks or gradient boosting trees, the need for explanations of their decisions has emerged. This demand has been driven by the increasing implementation of opaque models, in high-risk and critical areas like medicine, healthcare, industry, and law, which laid the foundation for modern research on explainable and interpretable artificial intelligence (XAI). Scientists’ efforts in designing XAI algorithms have been further supported by political initiatives such as DARPA’s XAI challenge [1], the European Union’s GDPR [2], and more recently, the EU AI Act [3].
@@ 1 Background & Summary
Our contribution addresses these gaps by presenting a dataset resulting from extensive user study on the comprehensibility of XAI algorithms across three distinct user groups with different information competencies. The dataset contains material collected from 39 participants during the interviews conducted by the information sciences research group. The participants were recruited from 149 candidates to form three groups that represented domain experts in the field of mycology (DE), students with a data science and visualization background (IT) and students from social sciences and humanities (SSH). Each group was provided with a set of explanations from a machine learning model trained to predict edible and non-edible mushrooms. The data used to train the ML model was in tabular format and is publicly available [6]. The machine learning model we used was Extreme Gradient Boosting (XGB) and explanations for its decision were prepared including state of the art model-agnostic algorithms such as SHAP, LIME, DICE and Anchor [7].
@@
Figure 1: Schema of a study, data collection and data format.
@@
We believe this dataset will significantly advance the field of XAI by providing the necessary tools to explore and improve the societal and comprehensibility aspects of AI systems. It will be a valuable resource for researchers aiming to develop more transparent and understandable AI technologies and foster research on the human-centered evaluation of XAI systems. The dataset is complemented with the source code, allowing one to reproduce the initial ML model and explanations.
2 Methods
2.1 Machine learning model and XAI algorithms
@@
The dataset used to train the classifier is a publically available dataset from the UC Irvine Machine Learning Repository [^1] See: https://archive.ics.uci.edu/dataset/848/secondary+mushroom+dataset. It contains data on 61,069 specimens from 173 mushroom species, categorized as either edible or inedible/poisonous. Specimens with unknown edibility were classified as inedible/poisonous. The dataset exclusively includes cap-and-stalk mushrooms with gill hymenophores and comprises both real observations and hypothetical data, the latter generated artificially from a smaller set of real-world mushroom observations.
@@
Regarding class distribution, the dataset is fairly balanced, with 33,888 instances labeled inedible or poisonous and 27,181 as edible, making approximately 55.49 This balance is crucial for accurate model training and performance evaluation, avoiding biases towards the majority class.
@@
The ML model we prepared and presented to the study participants uses the Gradient Boosting Classifier (XGBClassifier). To handle categorical variables, a one-hot encoder was used and missing data was addressed through imputation. Additionally, numerical variables were scaled to ensure uniformity in data handling. This preprocessing involved imputing missing values in numeric features with the median and scaling these features, while categorical variables were imputed with a placeholder value ’_NA_’. Consequently, the model operated on 82 features derived from the original 20 input variables. Renowned for its efficiency in classification tasks, the model exhibited an impressive accuracy rate of 99.97
@@
  • Rule-based explanations, such as Anchor and LUX,
@@
The explanations were compiled in a single PDF presentation with 14 slides, where each slide contains a visualization of the explanation generated with the selected XAI algorithm. There was also a 15th slide that indicated the end of the study. The sample of four slides depicting these visualizations is given in Figure 3. The type of slides and their ordering are presented in Figure 2. We encourage the reader to search for the full set of slides while reading the following description of methods used to craft the presentation available on Zenodo [9].
@@
Figure 3: Four samples slides with explanation visualizations used in the study including respectively: feature distributions, SHAP beeswarm, SHAP waterfall and LIME feature-importance plots.
2.1.1 Statistical descriptions and visualization of data
@@
2.1.2 Feature importance attribution-based explanations
@@
A SHAP "Waterfall" chart was utilized to interpret the model’s feature attributions for prediction. The custom title indicated the analysis of features that influence the prediction of the specified class, with the true class of the specimen also noted. This visualization delineated the contribution of each feature to the model’s prediction, with positive values indicating increased probability of the mushroom being classified as poisonous, and negative values the opposite.
The chart was accompanied by a legend where: $E[f(X)]$ signifies the baseline value or the mean model prediction, the color bands reflect the influence of each feature on the prediction of toxicity for the individual mushroom, and $f(x)$ is the final value equating to the model’s prediction for that specific mushroom. An explanatory note stated the waterfall chart’s function as displaying the impact of individual features on the toxicity prediction of a given mushroom. The term $E[f(X)]$ denoted the model’s average prediction across all observations, while $f(x)$ represents the prediction for the specific instance in question.
@@
2.1.3 Rule-based explanations
The study employed the ANCHOR package, a method of explainability for AI models that generates an "anchor", or a set of conditions that collectively influence the model’s classification of an instance with high precision.
An example observation matching the anchor criteria was displayed on the left side of the slide, representing the output from ANCHOR. The central part of the slide described the AI’s prediction process: for instances meeting all listed anchor conditions, the AI would predict the class as EDIBLE with 97.2 The anchor, composed of feature conjunctions that determine classification, is detailed in the upper right. Below it, the extent of influence these conditions had on the certainty of the classification was quantified. Since interactive features were not employed, the slide showed a non-interactive list of features. Replicating the checklist with fewer marked features offered insights into how different subsets contributed to the prediction, without providing full anchor conditions. Two slides were presented, one for the edible class and one for the inedible/poisonous class.
2.1.4 Counterfactual explanations
@@
2.2 Collection of empirical material
The collection of empirical material comprised two main phases: an online survey and individual qualitative interviews conducted using the think-aloud protocol (TAP) guidelines [10]. The survey involved non-intrusive questioning, while the TAP interviews required audio recording with explicit consent obtained from willing participants. Participation in both the survey and interviews was voluntary, and all collected data was anonymized. We opted for general professional vocabulary over industry-specific jargon to ensure clear communication without adding new content.
@@ 2.2 Collection of empirical material
Subsequently, these groups were assigned tasks and interviewed using advanced tools from social science fields such as TAP and MAXQDA for analysis. The interview part of the study included collection of qualitative material for further analysis with the TAP procedure. A total of 26 students (IT and SSH) and 13 experts took part in TAP. The IT group had 8 respondents from computer science, applied computer science, computer game science and electronic information processing. The SSH group therefore consisted of 18 respondents representing the following fields of study: psychology, practical psychology, applied psychology, information management, contemporary culture lab, journalism and social communication, and cognitive science. The SSH group was more extensive, as we were particularly interested in exploring the interaction experiences with XAI techniques of respondents not professionally involved in data analysis and visualisation in the context of artificial intelligence and computer science.
Regarding the tasks performed by the respondents during TAP, we utilized a list of three main tasks.
In the Task 1 participants were asked to review a list of questions related to each explanation and answer as many as they could. There were 17 questions in total, each corresponding to a specific explanation type. Additionally, there were three problem-solving questions where participants were asked to classify mushrooms based on the knowledge they gained from the explanations. In the second task (Task 2), participants were asked to arrange the explanations in the order they found most useful, removing any unnecessary explanations and possibly adding general ideas on what explanations should be included and their preferred order. In the third task (Task 3), participants were asked a general question: "What improvements can be made to the prepared materials to make them more useful to you?"
@@
The TAP study was followed by a thematic analysis which is is a qualitative data analysis approach that emphasizes identifying and describing prominent themes within the data. It also typically involves comparing themes across different conditions or populations and exploring the relationships between these themes [12, 13, 14]. Thematic analysis comprise of six consecutive stages. The first stage involved an exploratory phase for data reconnaissance and familiarization. The second phase consisted of the independent, inductive generation of codes by two researchers. Next, themes were created and verified, during which the initial and subsequent codebooks were developed by deductive referencing codes to the primary research goals of our research which was focused on evaluation of comprehensibility of XAI. This stage includes refinement of original codes and themes once the analysis progress. Finally, the analysis concluded with the compilation of results, highlighting the co-occurrences of themes among different observations. [15].
Between January 17, 2024, and March 4, 2024, a total of 66 items were collected for coding, including respectively: 13 transcripts of research video sessions with domain experts, as well as 27 transcripts of study conducted with students and 26 notes prepared by the think-aloud protocol facilitators after the study’s conclusion. MAXQDA 2024 Analytics Pro software (the UJ licence is valid until 5 January 2025) was used for coding the materials and their initial analysis and visualisation, as was Tableau Software for visualising the results. The dataset presented in our work contains both the codebook designed during the analysis as well as the final result of the analysis in a form of codes’ occurences among different participants.
3 Data Records
The general structure of the dataset is described in Table 1. The files that contain in their names [RR]_[SS]_[NN] contain the individual results obtained from particular participant. The meaning of the prefix is as follows:
@@ 3 Data Records
QUESTIONS.csv List of questions asked to participants during interviews. Polish
@@
PROBLEMS_RESPONSES.csv Responses to the problems listed in PROBLEMS.csv for each participant. Polish
@@
Figure 4: Schema of data records. Arrows represent intended way of joining tables, while colors denote keys in different tables that can be used to join tables in other ways.
3.1 Initial survey results
The initial survey dataset contains self-assessment questions that all participant candidates were required to answer. The survey comprised 18 questions, 12 of which pertained directly to self-assessing knowledge and skills in mushrooming and data visualisation, as well as acquiring information about their origin and experience. The survey included a question about the field of study to ensure appropriate representation of students pursuing humanities, social sciences, and computer science. Five questions in the questionnaire directly related to giving consent for data use and participation in further research, as well as confirming familiarity with the GDPR clause. The questionnaire aimed to gather information and assess the knowledge, qualifications, experience, and competences of the participants. It consisted of six open-ended questions and six closed questions. The latter included an ’other’ option, allowing participants to provide their own answer. These questions enable the grouping of participants based on their domain knowledge and data analysis literacy, which is crucial for selecting the final set of participants and categorizing them into DE, IT, and SSH groups. The description of a content of this set of data is given in Table 2. The survey table contains 154 records, including 4 candidates used in the pilot study who were not included in the final dataset. The goal of the pilot study was to identify potential issues, refine research methodologies, and ensure the feasibility of the study design; therefore, the participants in this stage were not considered valid subjects for the final dataset.
@@ 3.1 Initial survey results
participant_id Unique identifier of a participant who took part in the interviews. This column should be used to join records from other tables.
@@
3.2 Labelled interview transcripts
@@
The key column to identify the stage of the interview is slide_id column. Besides the ID of a currently analyzed slide, there are also three other special IDs that mark different parts of the interview. The interpretation of the special IDs is as folows: __S00__ indicates the begining of the core part of the interview with slide analysis, before that the description of the study is introduced to the participant; __S99__ represent the beginning of the section where the participants analyze visualization order, the slides ids are not assigned in this section, due to dynamic slide switching by participants; __S15__ represents the end of the slide analysis section; __S88__ represents the beginning of the problem solving section.
@@
3.3 Explanation-guided mushroom classification by users
The participants were asked to solve three problems related to mushroom classification as edible and non-edible based on the explanations and other material they were provided. All of the participants were asked about the same three instances that contain two edible and one non-edible mushroom with varying ML model classification probability assigned to each instance. The table containing this instances along with ML model predictions is described in detail in Table 6, while the responses of the participants were stored in a table of a schema defined by Table 7. The participants were also asked to indicate their level of certainty of prediction that can later be confronted with ML model classification probability. The DE group of participants was not solving this tasks, hence the table is missing participant IDs related to domain experts in mycology.
@@ 3.3 Explanation-guided mushroom classification by users
3.4 Modification of explanation visualization
One of the tasks assigned to participants was to suggest modifications to the provided explanations to improve overall user experience and enhance their usability. Participants could reorder the explanations, remove selected ones, or add custom slides with additional explanations or information they deemed useful. The schema of the table containing records from this part of the study is shown in Table 8. The DE group of participants was not solving this tasks neither.
@@ 3.4 Modification of explanation visualization
3.5 Results from thematic analysis with MAXQDA
@@
4 Technical Validation
4.1 Classification model training and validation
The creation of a ML model was followed by the standard practices in data science including numerical data scaling, categorical variables encoding and missing values imputation. The detest was balanced, so not oversampling nor under-sampling techniques were applied that could alter the original data distribution. The model training was performed on train set which comprises 70 The hyper-parameter tuning was performed with grid search and five-fold cross validation. With that approach, we achieved the performance of the model at the level of 99,97 For the full transparency and reproducibility of this procedure, we provide a source code of the model training and data preprocessing [^2] See: https://gitlab.geist.re/pro/xai-fungi..
4.2 Study design and data acquisition
@@
4.3 Data post-processing
The post-processing of the results, involving the transformation of all gathered material into structured CSV format as presented in previous sections, underwent subsequent verification steps. Initially, the technical integrity of the material was assessed by two researchers to ensure no data records were missing from the original sources. Following this, the transcripts and their manual labeling were reviewed by three independent researchers to confirm the accuracy and integrity of this stage. Finally, we performed technical analysis of the dataset, ensuring that all of the tables contain valid IDs and can be merged according to schema depicted in Figure 4.
@@ 4.3 Data post-processing
In this example in Listing 1, we present how to read sample records from a dataset, how to link the records with other tables and trigger basic analysis with scikit-learn library [22]. The example assumes that all of the files are in the zenodo folder and the transcripts were unzipped into zenodo/transripts folder.
import pandas as pd
import seaborn as sns
#select participant which ouy would like to investigate
participant_id = ’MW_SSH_05’
problems = pd.read_csv(’./zenodo/PROBLEMS.csv’)
responses = pd.read_csv(’./zenodo/PROBLEMS_RESPONSES.csv’)
merged_problems = pd.merge(responses,problems, on=’problem_id’)
prediction_features = [’participant_decision’,’participant_certaintity’,
’model_class’,’model_probability’]
print(merged_problems[merged_problems[’participant_id’]==participant_id][prediction_features])
#retrieve related text that correspond to the problem P1 of a particiapnt MZ_SSH_04
transcripts = pd.read_csv(f’./zenodo/transcripts/{participant_id}.csv’)
transcripts[’problem_id’] = transcripts[’problem_id’].ffill().dropna()
transcripts[transcripts[’problem_id’] == ’__P1__’].dropna(axis=1)
Listing 1: Merging two tables representing user responses to classification problems and retrieving related text from transcript.
The code reads two tables containing problems and problem responses, merges them into one table, retrieves records for participants encoded as MW_SSH_05, and prints the participant responses along with transcript fragments related to the interview segment where the participant solved the problem.

### v2

# Dataset resulting from the user study on comprehensibility of explainable AI algorithms

@@ Abstract
Background & Summary
With the rapid development of black-box machine learning (ML) models, such as deep neural networks or gradient boosting trees, the need for explanations of their decisions has emerged. This demand has been driven by the increasing implementation of opaque models, in high-risk and critical areas like medicine, healthcare, industry, and law, which laid the foundation for modern research on explainable and interpretable artificial intelligence (XAI). Scientists’ efforts in designing XAI algorithms have been further supported by political initiatives such as Defense Advanced Research Projects Agency’s (DARPA) XAI challenge [1], the European Union’s (EU) General Data Protection Regulation (GDPR) [2], and more recently, the EU AI Act [3].
@@
Our contribution addresses these gaps by presenting a dataset resulting from extensive user study on the comprehensibility of XAI algorithms across three distinct user groups with different information competencies. The dataset contains material collected from 39 participants during the interviews conducted by the information sciences research group. The final 39 participants were recruited from 149 candidates to form three groups that represented domain experts in the field of mycology (DE), students with a data science and visualization background (IT) and students from social sciences and humanities (SSH). The IT group consisted of 8 participants, the SSH group included 18 participants, and the DE group comprised 13 domain experts. Each group was provided with a set of explanations from a machine learning model trained to predict edible and non-edible mushrooms.
The data used to train the ML model was in tabular format and is publicly available [6]. The machine learning model we used was Extreme Gradient Boosting (XGB) and explanations for its decision were prepared including state of the art model-agnostic algorithms such as SHapley Additive exPlanations (SHAP), Local Interpretable Model-agnostic Explanations (LIME), Diverse Counterfactual Explanations (DICE) and Anchor [7].
@@
Figure 1: Schema of a study, data collection and data format. The columns which contain self-explanatory headings are put in square brackets in the figure, indicating that there is a multiple columns related to some type of information, e.g. survey questions in SURVEY or feature names in PROBLEMS.
@@
This dataset addresses the gap identified in theoretical works, such as Miller’s [5], which emphasize the need for empirical evaluations of XAI comprehensibility but lack real-world user studies. Our dataset provides a richer and more comprehensive empirical foundation by including a diverse set of user interviews, enabling a deeper understanding of how various user groups—such as domain experts, data science students, and social science students—comprehend XAI algorithms. Unlike studies such as [8, 9, 10], which focus on a specific professional group, or evaluate XAI methods within a single user group, this dataset offers a broader evaluation by examining a wider spectrum of user expertise, thus providing more comprehensive insights into the challenges XAI faces across different domains.
Our goal was to provide a dataset from the study that can serve as a resource to replicate our experiment but also for in depth analysis of the empirical material we collected. We envision broader applications of our dataset, such as its use in multimodal explainability methods, where interview transcripts and XAI visualizations could aid in developing and training multimodal (text-image) conversational agents for interactive XAI. We believe our dataset has the potential to significantly advance XAI research, extending far beyond the examples outlined above. In addition, our dataset offers valuable opportunities for personalization and user modeling in XAI. The transcripts capture detailed user interactions and preferences, enabling the development of user models that reflect differences in domain knowledge, explanation needs, and cognitive strategies. These models can inform the design of adaptive XAI systems that tailor explanations to individual users, enhancing comprehension and trust. This makes the dataset a strong foundation for building preference-aware recommender systems and interactive interfaces that support personalized, human-centric explainability. The dataset is complemented with the source code, allowing one to reproduce the initial ML model and explanations.
Methods
Machine learning model and XAI algorithms
@@ Methods
The dataset used to train the classifier is a publicly available dataset from the University of California Irvine Machine Learning Repository [11]. It contains data on 61,069 specimens from 173 mushroom species, categorized as either edible or inedible/poisonous. Specimens with unknown edibility were classified as inedible/poisonous. The dataset exclusively includes cap-and-stalk mushrooms with gill hymenophores and comprises both real observations and hypothetical data, the latter generated artificially from a smaller set of real-world mushroom observations. The hypothetical and artificially generated data was created in accordance to the description of the morphological structure of the mushrooms obtained from the Mushrooms and Toadstools book  [12].
@@
Regarding class distribution, the dataset is fairly balanced, with 33,888 instances labeled inedible or poisonous and 27,181 as edible, making approximately 55.49% of the entries inedible/poisonous. This balance is crucial for accurate model training and performance evaluation, avoiding biases towards the majority class.
@@
The ML model we prepared and presented to the study participants uses the XGB classifier. The motivation for XGB was its excellent performance on the dataset. To handle categorical variables, a one-hot encoder was used and missing data was addressed through imputation. Additionally, numerical variables were scaled to ensure uniformity in data handling. This preprocessing involved imputing missing values in numeric features with the median and scaling these features, while categorical variables were imputed with a placeholder value ’_NA_’. Consequently, the model operated on 82 features derived from the original 20 input variables. Renowned for its efficiency in classification tasks, the model exhibited an impressive accuracy rate of 99.97%.
@@
  • Rule-based explanations, such as Anchor and Local Universal Explainer (LUX) [13],
@@
The selection of explainability methods - Anchor, DICE, LIME, and SHAP - was driven by the need to cover a diverse range of well-established explainers in the field. We chose the most popular and widely recognized methods, each representing a distinct type of explanation (e.g., local, feature-based, rule-based) to ensure comprehensive coverage.
@@
The explanations were compiled in a single PDF presentation with 14 slides, where each slide contains a visualization of the explanation generated with the selected XAI algorithm. There was also a 15th slide that indicated the end of the study. The type of slides and their ordering are presented in Figure 2. We encourage the reader to search for the full set of slides while reading the following description of methods used to craft the presentation available on Zenodo [15].
@@
Statistical descriptions and visualization of data
@@
Feature importance attribution-based explanations
@@
A SHAP "Waterfall" chart was utilized to interpret the model’s feature attributions for prediction. The custom title indicated the analysis of features that influence the prediction of the specified class, with the true class of the specimen also noted. This visualization delineated the contribution of each feature to the model’s prediction, with positive values indicating increased probability of the mushroom being classified as poisonous, and negative values the opposite. The chart was accompanied by a legend where: $E[f(X)]$ signifies the baseline value or the mean model prediction, the color bands reflect the influence of each feature on the prediction of toxicity for the individual mushroom, and $f(x)$ is the final value equating to the model’s prediction for that specific mushroom. An explanatory note stated the waterfall chart’s function as displaying the impact of individual features on the toxicity prediction of a given mushroom. The term $E[f(X)]$ denoted the model’s average prediction across all observations, while $f(x)$ represents the prediction for the specific instance in question.
@@
Rule-based explanations
The study employed the ANCHOR package, a method of explainability for AI models that generates an "anchor", or a set of conditions that collectively influence the model’s classification of an instance with high precision. An example observation matching the anchor criteria was displayed on the left side of the slide, representing the output from Anchor. The central part of the slide described the AI’s prediction process: for instances meeting all listed anchor conditions, the AI would predict the class as EDIBLE with 97.2% certainty. The anchor, composed of feature conjunctions that determine classification, is detailed in the upper right. Below it, the extent of influence these conditions had on the certainty of the classification was quantified. Since interactive features were not employed, the slide showed a non-interactive list of features. Replicating the checklist with fewer marked features offered insights into how different subsets contributed to the prediction, without providing full anchor conditions. Two slides were presented, one for the edible class and one for the inedible/poisonous class.
Counterfactual explanations
@@
Collection of empirical material
The collection of empirical material comprised two main phases: an online survey and individual qualitative interviews conducted using the think-aloud protocol (TAP) guidelines [16]. The survey involved non-intrusive questioning, while the TAP interviews required audio recording with explicit consent obtained from willing participants. Participation in both the survey and interviews was voluntary, and all collected data was anonymized. Before starting the survey, participants were asked to confirm that their participation was voluntary and that they acknowledged the data collected during the experiment would be shared for scientific purposes, in accordance to GDPR regulations. Participants were also asked for their consent to make all empirical material - including interview transcripts and other data collected during the study - available for publication after anonymization. We opted for general professional vocabulary over industry-specific jargon to ensure clear communication without adding new content.
@@
Subsequently, these groups were assigned tasks and interviewed using advanced tools from social science fields such as TAP and MAXimum Qualitative Data Analysis (MAXQDA) for analysis. The interview part of the study included collection of qualitative material for further analysis with the TAP procedure. The MAXQDA methodology was chosen due to its widespread use in qualitative and mixed-methods research, particularly in user studies, for systematically analyzing and interpreting textual and multimedia data. It also enables the integration of qualitative insights with quantitative data, supporting the development of more user-centered and interactive explainability methods.
In the TAP study, we employed thematic analysis to systematically explore user perspectives, identify recurring themes, and uncover meaningful insights that quantitative methods alone might overlook. This approach is especially well-suited for capturing the depth and complexity of user experiences, which is crucial for understanding how users interact with AI explanations.
A total of 26 students (IT and SSH) and 13 experts took part in TAP. The IT group had 8 respondents from computer science, applied computer science, computer game science and electronic information processing. The SSH group therefore consisted of 18 respondents representing the following fields of study: psychology, practical psychology, applied psychology, information management, contemporary culture lab, journalism and social communication, and cognitive science. The SSH group was more extensive, as we were particularly interested in exploring the interaction experiences with XAI techniques of respondents not professionally involved in data analysis and visualisation in the context of artificial intelligence and computer science.
Regarding the tasks performed by the respondents during TAP, we utilized a list of three main tasks. In the Task 1 participants were asked to review a list of questions related to each explanation and answer as many as they could. There were 17 questions in total, each corresponding to a specific explanation type. Additionally, there were three problem-solving questions where participants were asked to classify mushrooms based on the knowledge they gained from the explanations. In the second task (Task 2), participants were asked to arrange the explanations in the order they found most useful, removing any unnecessary explanations and possibly adding general ideas on what explanations should be included and their preferred order. In the third task (Task 3), participants were asked a general question: "What improvements can be made to the prepared materials to make them more useful to you?"
@@
The TAP study was followed by a thematic analysis which is a qualitative data analysis approach that emphasizes identifying and describing prominent themes within the data. It also typically involves comparing themes across different conditions or populations and exploring the relationships between these themes [18, 19, 20]. Thematic analysis comprises of six consecutive stages. The first stage involved an exploratory phase for data reconnaissance and familiarization. The second phase consisted of the independent, inductive generation of codes by two researchers. Next, themes were created and verified, during which the initial and subsequent codebooks were developed by deductive referencing codes to the primary research goals of our research which was focused on evaluation of comprehensibility of XAI. This stage includes refinement of original codes and themes once the analysis progress. Finally, the analysis concluded with the compilation of results, highlighting the co-occurrences of themes among different observations. [21].
Between January 17, 2024, and March 4, 2024, a total of 66 items were collected for coding, including respectively: 13 transcripts of research video sessions with domain experts, as well as 27 transcripts of study conducted with students and 26 notes prepared by the think-aloud protocol facilitators after the study’s conclusion. MAXQDA 2024 Analytics Pro software (the UJ licence is valid until 5 January 2025) was used for coding the materials and their initial analysis and visualisation, as was Tableau Software for visualising the results. The dataset presented in our work contains both the codebook designed during the analysis as well as the final result of the analysis in a form of codes’ occurrences among different participants.
Data Records
The complete dataset was deposited in the Zenodo repository [15]. The general structure of the dataset is described in Table 1. The files that contain in their names [RR]_[SS]_[NN] contain the individual results obtained from particular participant. The meaning of the prefix is as follows:
@@
SURVEY_EN.csv Translation of the SURVEY content to English. English
@@
QUESTIONS.csv List of questions asked to participants during interviews. Polish and English
@@
PROBLEMS_EN.csv Translation of the PROBLEMS content to English English
PROBLEMS_RESPONSES.csv Responses to the problems listed in PROBLEMS.csv for each participant. Polish and English
@@
ORIGINAL_VISUALIZATIONS_EN.pdf PDF file containing original visualizations presented to participants translated into English. English
@@
Figure 3: Scheme of data records. Arrows represent intended way of joining tables, while colors denote keys in different tables that can be used to join tables in other ways. Tables containing translations should be used for preview only to better understand the content. Column name between square brackets mean that it actually represents a set of columns which names are self-explanatory and there were not included in the figure for the sake of readability.
Initial survey results
The initial survey dataset contains self-assessment questions that all participant candidates were required to answer. The survey comprised 18 questions, 12 of which pertained directly to self-assessing knowledge and skills in mushrooming and data visualisation, as well as acquiring information about their origin and experience. The survey included a question about the field of study to ensure appropriate representation of students pursuing humanities, social sciences, and computer science. Five questions in the questionnaire directly related to giving consent for data use and participation in further research, as well as confirming familiarity with the GDPR clause. The questionnaire aimed to gather information and assess the knowledge, qualifications, experience, and competences of the participants. It consisted of six open-ended questions and six closed questions. The latter included an ’other’ option, allowing participants to provide their own answer. These questions enable the grouping of participants based on their domain knowledge and data analysis literacy, which is crucial for selecting the final set of participants and categorizing them into DE, IT, and SSH groups. The description of a content of this set of data is given in Table 2. The survey table contains 154 records, including 4 candidates used in the pilot study who were not included in the final dataset. The goal of the pilot study was to identify potential issues, refine research methodologies, and ensure the feasibility of the study design; therefore, the participants in this stage were not considered valid subjects for the final dataset. The missing values in this table indicate that the candidate did not provide an answer. Candidates who were not selected for interviews do not have a participant_id assigned.
@@
participant_id Unique identifier of a participant who took part in the interviews. This column should be used to join records from other tables. Candidates with missing ID were not selected for the interviews.
comment The additional comment related to the entry.
@@
Labelled interview transcripts
@@
The key column to identify the stage of the interview is slide_id column. Besides the ID of a currently analyzed slide, there are also three other special IDs that mark different parts of the interview. The interpretation of the special IDs is as follows: __S00__ indicates the beginning of the core part of the interview with slide analysis, before that the description of the study is introduced to the participant; __S99__ represent the beginning of the section where the participants analyze visualization order, the slides ids are not assigned in this section, due to dynamic slide switching by participants; __S15__ represents the end of the slide analysis section; __S88__ represents the beginning of the problem solving section.
@@
Explanation-guided mushroom classification by users
The participants were asked to solve three problems related to mushroom classification as edible and non-edible based on the explanations and other material they were provided. All of the participants were asked about the same three instances that contain two edible and one non-edible mushroom with varying ML model classification probability assigned to each instance. The table containing this instances along with ML model predictions is described in detail in Table 6, while the responses of the participants were stored in a table of a schema defined by Table 7. The participants were also asked to indicate their level of certainty of prediction that can later be confronted with ML model classification probability. Missing values in this table indicates that the participant did not provide an answer. The DE group of participants was not solving this tasks, hence the table is missing participant IDs related to domain experts in mycology.
@@
prediction_decision_en Translation to English of the class assigned by the participant to the given problem.
@@
prediction_certainty_en Translation to English of the certainty of the participant’s decision.
@@
Modification of explanation visualization
One of the tasks assigned to participants was to suggest modifications to the provided explanations to improve overall user experience and enhance their usability. Participants could reorder the explanations, remove selected ones, or add custom slides with additional explanations or information they deemed useful. The schema of the table containing records from this part of the study is shown in Table 8. The DE group of participants was not solving this tasks neither. Missing values in the slide_id or original_order column indicate that the slide was not part of the original presentation and was suggested for addition by the participant. Missing values in the new_order column indicate that the slide was recommended for removal.
@@
Results from thematic analysis with MAXQDA
@@
Technical Validation
Classification model training and validation
The creation of a ML model was followed by the standard practices in data science including numerical data scaling, categorical variables encoding and missing values imputation. The dataset was balanced, so not oversampling nor under-sampling techniques were applied that could alter the original data distribution. The model training was performed on train set which comprises 70% of all samples randomly drawn for the full dataset. The hyper-parameter tuning was performed with grid search and five-fold cross validation and tested on the test-set not used for any training task before, to mitigate the overfitting risks. With this approach, we achieved a model performance of 99.97% accuracy, which is consistent with results reported in other studies using this dataset [11] For the full transparency and reproducibility of this procedure, we provide a source code of the model training and data preprocessing (see: https://gitlab.geist.re/pro/xai-fungi).
Study design and data acquisition
@@
Data post-processing
The post-processing of the results, involving the transformation of all gathered material into structured CSV format as presented in previous sections, underwent subsequent verification steps. Initially, the technical integrity of the material was assessed by two researchers to ensure no data records were missing from the original sources. Following this, the transcripts and their manual labeling were reviewed by three independent researchers to confirm the accuracy and integrity of this stage. Finally, we performed technical analysis of the dataset, ensuring that all of the tables contain valid IDs and can be merged according to schema depicted in Figure 3. The part of the dataset that were translated into English, were translated with a help of ChatGPT, and verified by us manually.
@@
In the online repository (see https://gitlab.geist.re/pro/xai-fungi), we present how to read sample records from a dataset, how to link the records with other tables and trigger basic analysis with scikit-learn library [28]. The examples assume that all of the files are in the zenodo folder and the transcripts were unzipped into zenodo/transripts folder.